Will is 4 years older than Phil. Three years from now, Will will be twice as old as Phil was last year. Find the present ages.
Found 2 solutions by stanbon, davethejackal:
Answer by stanbon(75887)

Will is 4 years older than Phil. Three years from now, Will will be twice as old as Phil was last year. Find the present ages.
--------
Equations:
w = p + 4
w-3 = 2(p-1)
--------------
Substitute and solve for "w":
p+4-3 = 2p-2
p+1 = 2p-2
p = 3 yrs (Phil's age now)
---------
Solve for "w":
w = p+4
w = 3+4
w = 7 (Will's age now)
------
Cheers,
Stan H.
==================
------------------

Answer by davethejackal(28)

Will is 4 years older than Phil. Three years from now, Will will be twice as old as Phil was last year. Find the present ages.

Let w= Will's age, p=Phil's age

[1] Will is 4 years older than Phil:
w = p + 4

[2] Three years from now, Will will be twice as old as Phil was last year:
w+3 = 2(p-1)
therefore w = 2(p-1)-3 = 2p-5

Combining [1] and [2]
p+4 = 2p-5 => p=9

since w=p+4 =9+4 = 13.

Will is 13, Phil is 9.